### v1

# A coisotropic embedding theorem for pre-multisymplectic manifolds

A coisotropic embedding theorem for pre-multisymplectic manifolds
L. Schiavone1,2 [Uncaptioned image]
1Dipartimento di Matematica e Applicazioni Renato Caccioppoli, Università degli Studi di Napoli Federico II
2
e-mail: luca.schiavone@unina.it
Acknowledgements
I am indebted to prof. A. Ibort for making me aware of the existence of the coisotropic embedding theorem and its potential applications as well as for the huge amount of inspiring discussions of the last years. I am also indebted to prof. G. Marmo for encouraging me to think of alternative ways to depict the geometry behind Gotay’s coisotropic embedding theorem. I also want to thank prof. A. Zampini, F.M. Ciaglia and F. Di Cosmo for many useful discussions. I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006. I am a member of the GNSGA (Indam).
1 Introduction
Multisymplectic geometry has emerged as a fundamental framework for the geometric description of Field Theories, extending the role that symplectic geometry plays in Classical Mechanics, as testified by the huge, even if not exhaustive, list of contributions [GPR69, GPR71, Kij73, GS73, KS76, TK79, AA80, Got91a, Got91b, BSF88, GMS97, CCI91, FF03, EEMLRR96, EEMLRR00, RR09, Kru15, CDI+20b, CDI+20a, CDI+24, CDI+22b, CDI+22c] (see also references therein). In this setting, the tangent and cotangent bundles of a configuration space are replaced by the first-order jet bundle of a fibration and its dual, providing a natural arena for analyzing the equations of motion and symmetries of Field Theories. These jet bundles are examples of pre-multisymplectic manifolds, which generalize the notion of pre-symplectic geometry to higher-degree forms. In this framework, to any Lagrangian (or Hamiltonian), a pre-multisymplectic structure is associated, that allows for an intrinsic formulation of a variational principle providing the equations of motion of the theory.
In previous contributions [CDI+20b, CDI+20a, CDI+24, CDI+22b, CDI+22c], we showed how, at least locally close to some Cauchy hypersurface, such equations of motion can be formulated in terms of (infinite-dimensional) pre-symplectic Hamiltonian systems. In particular, we established that the space of solutions of the equations of motion inherits a pre-symplectic structure. The kernel of this structure descends from the kernel of the pre-multisymplectic form mentioned above and encodes the gauge symmetries of the theory.
To define a Poisson structure on the space of solutions, we employed a regularization technique based on M. Gotay’s coisotropic embedding theorem [Got82] (see also [GS90] for an equivariant version of the theorem and [OP05] for a more modern approach). Gotay’s theorem provides a powerful tool in symplectic geometry, ensuring that any pre-symplectic manifold can be embedded as a coisotropic submanifold into a symplectic manifold, referred to as its symplectic thickening. In [CDI+24, CDI+22b, CDI+22c], we used Gotay’s theorem to coisotropically embed the space of solutions of the equations of motion of Classical Field Theories into a symplectic manifold where a Poisson structure is canonically defined. We then studied whether and how this Poisson structure projects back to a Poisson structure on the space of solutions.
Having in mind the previous statement that Field Theories can be locally formulated in terms of (infinite-dimensional) pre-symplectic Hamiltonian systems, the coisotropic embedding theorem has been used by the author also to provide a one-to-one correspondence between symmetries and constants of the motion [CDI+22a] and to provide a solution for the inverse problem of the calculus of variations for a class of implicit differential equations [Sch24b, Sch24a].
While effective in all the cases mentioned above, this regularization approach requires handling infinite-dimensional spaces, as the space of solutions is typically infinite-dimensional. To circumvent the technical challenges associated with infinite-dimensional spaces, it may be useful to develop a multisymplectic analogue of Gotay’s theorem, allowing coisotropic regularization to be carried out directly at the finite-dimensional level of the underlying pre-multisymplectic manifold.
The aim of this manuscript is to prove a coisotropic embedding theorem for pre-multisymplectic manifolds. Specifically, we will show that any pre-multisymplectic manifold can be embedded as a coisotropic submanifold into a larger multisymplectic manifold. This generalization provides a finite-dimensional framework for coisotropic regularization and opens new avenues for the study of multisymplectic geometry and its applications to Field Theories.
It is worth pointing out that recent work has focused on the interplay between coisotropic submanifolds and multisymplectic geometry, particularly in the opposite direction taken in this manuscript, namely that of multisymplectic reduction. In particular, in [de 24] the authors analyze coisotropic submanifolds of pre-multisymplectic manifolds and study conditions under which they inherit multisymplectic structures.
The structure of this paper is as follows. In Section 2, we recall preliminary notions about pre-multisymplectic manifolds that will be used throughout the manuscript. Section 3 is devoted to the statement and proof of the main theorem.
2 Preliminaries
Definition 2.1 (Multisymplectic manifold).
A $k$-multisymplectic manifold is a smooth differential manifold $\mathcal{M}$ equipped with a closed and non-degenerate differential $k$-form $\omega$.
Definition 2.2 (Pre-multisymplectic manifold).
A $k$-pre-multisymplectic manifold is a smooth differential manifold $\mathcal{M}$ equipped with a closed differential $k$-form $\omega$.
Definition 2.3 ($\ell$-multisymplectic orthogonal).
Given a multisymplectic manifold $(\mathcal{M},\,\omega)$ and a submanifold $\mathcal{N}\subset\mathcal{M}$, the $\ell$-multisymplectic orthogonal of $\mathcal{N}$ in $\mathcal{M}$ at $n\in\mathcal{N}$ is
[…]
Evidently
$\mathbf{T}_{n}\mathcal{N}^{\perp,\,j}\subseteq\mathbf{T}_{n}\mathcal{N}^{\perp,\,i}\,,\;\;\;\text{if }\,\,i>j\,,$ (2)
and
$\mathbf{T}_{n}\mathcal{N}^{\perp,\,\ell}\,=\,\mathbf{T}_{n}\mathcal{M}\,,\;\;\;\text{if }\,\,\ell>k\,.$ (3)
Definition 2.4 ($\ell$-Coisotropic submanifold).
A submanifold $\mathcal{N}$ of a multisymplectic manifold $(\mathcal{M},\,\omega)$ is $\ell$-coisotropic if $\mathbf{T}_{n}\mathcal{N}^{\perp,\,\ell}\subseteq\mathbf{T}_{n}\mathcal{N}\,,\;\;\;\forall\,\,n\in\mathcal{N}$.
Definition 2.5 (Bundle of $k$-forms).
Given a smooth differential manifold $\mathcal{M}$, we will denote by $\Lambda^{k}(\mathcal{M})$ the bundle of differential $k$-forms on $\mathcal{M}$, namely the vector bundle over $\mathcal{M}$ whose typical fibre at $m\in\mathcal{M}$ is $\bigwedge^{k}\mathbf{T}^{\star}_{m}\mathcal{M}$. We will denote by $\pi$ the canonical projection $\pi\;:\;\;\Lambda^{k}(\mathcal{M})\to\mathcal{M}$.
Given a smooth $d$-dimensional differential manifold $\mathcal{M}$, we will usually denote by
$\left\{\,q^{j}\,\right\}_{j=1,...,d}$ (4)
a system of local coordinates on it and by
$\left\{\,q^{j},\,p_{j_{1}...j_{k}}\,\right\}_{j,j_{1},...,j_{k}=1,...,d}$ (5)
an adapted system of local coordinates on $\Lambda^{k}(\mathcal{M})$.
Consider a $k$-pre-multisymplectic manifold $(\mathcal{M},\,\omega)$. The kernel of $\omega$ is, at each $m\in\mathcal{M}$, a subspace of $\mathbf{T}_{m}\mathcal{M}$, denoted by $K_{m}$. We will always assume $\omega$ to have constant rank, so that $K_{m}$ and $K_{n}$ are isomorphic for any pair $m,\,n\in\mathcal{M}$. Moreover, we will also always assume that $K_{m}$ provides a completely integrable distribution on $\mathcal{M}$ [^1] Note that the complete integrability of the distribution generated by $K_{m}$ is ensured by the closure of $\omega$ only when $k=2$ by Frobenius theorem [AMR88]. so that there exists a foliation of $\mathcal{M}$ such that the tangent space to each leaf at each point $m\in\mathcal{M}$ coincides with $K_{m}$ and there exists a unique leaf passing through any point of $\mathcal{M}$. The space of leaves of the foliation is the space of equivalence classes of points on $\mathcal{M}$ belonging to the same leaf and will be denoted by
$\mathcal{M}/\sim_{K}\,.$ (6)
Sometimes, we will consider on $\mathcal{M}$ a system of local coordinates adapted to such a foliation
$\left\{\,q^{j}\,\right\}_{j=1,...,d}\,=\,\left\{\,x^{a},\,f^{A}\,\right\}_{a=1,...,l;A=1,...,r}\,,$ (7)
where $l$ is the rank of $\omega$ and $r$ is the dimension of its kernel. The coordinates $x^{a}$ are a system of coordinates on the space of leaves of the foliation, namely, locally each leaf is a level set of the type
$\mathcal{F}_{c}\,=\,\left\{\,x^{a}\,=\,c^{a}\,,\;\;\;c^{a}\in\mathbb{R}\;\;\;\forall\,\,a=1,...,l\,\right\}\,,$ (8)
whereas the coordinates $f^{A}$ (for fixed values of the $x^{a}$’s) individuates a point on each leaf. In this system of coordinates, $K_{m}$ reads
$K_{m}\,=\,\langle\left\{\,\frac{\partial}{\partial f^{A}}\,\right\}_{A=1,...,r}\rangle\,.$ (9)
A complement to $K_{m}$ at each $m\in\mathcal{M}$ is not canonically defined. Indeed, each choice of a complement $W_{m}$ such that
$\mathbf{T}_{m}\mathcal{M}\,=\,K_{m}\oplus W_{m}\,,$ (10)
amounts to the choice of a connection on $\mathcal{M}$.
Definition 2.6 (Connection on $\mathcal{M}$).
A connection on $\mathcal{M}$ is an idempotent smooth (1, 1)11(1,\,1)( 1 , 1 )-tensor field on $\mathcal{M}$ whose image, at each $m\in\mathcal{M}$, is $K_{m}$.
Locally a connection can be written as
$P\,=\,({\rm d}f^{A}-P^{A}_{a}{\rm d}x^{a})\otimes\frac{\partial}{\partial f^{A}}\,,$ (11)
where the functions $P^{A}_{a}$ are the so-called connection coefficients (see, for instance [GMS10, App. $14$] or [GMS09, Sec. $1.3$]). By means of $P$, the tangent space to $\mathcal{M}$ at each point splits into the direct sum of tangent vectors in the image of $P$ and tangent vectors in the kernel of $P$. The image of $P$ reads
$K_{m}\,=\,\langle\left\{\,\frac{\partial}{\partial f^{A}}\,\right\}_{A=1,...,r}\rangle\,,$ (12)
and is usually referred to as the space of vertical vectors, whereas the kernel reads
$W_{m}\,=\,\langle\left\{\,W_{a}\,:=\,\frac{\partial}{\partial x^{a}}+P^{A}_{a}\frac{\partial}{\partial f^{A}}\,\right\}_{a=1,...,l;A=1,...,r}\rangle\,,$ (13)
and is usually referred to as the space of horizontal vectors. It is easy to show that under a change of local chart, the point-wise splitting into vertical and horizontal tangent vectors is preserved and, thus, the tangent bundle to $\mathcal{M}$ split as the direct sum
$\mathbf{T}\mathcal{M}\,=\,\mathbf{V}^{P}\oplus\mathbf{H}^{P}\,,$ (14)
where $\mathbf{V}^{P}$ (resp. $\mathbf{H}^{P}$) is the vector subbundle of $\mathbf{T}\mathcal{M}$ whose sections are, at each point, the vertical (resp. horizontal) vectors defined by $P$.
[…]
$\mathbf{T}\mathcal{M}\,=\,\mathbf{V}^{R}\oplus\mathbf{H}^{R}\,,$ (15)
the splitting of $\mathbf{T}\mathcal{M}$ induced by the connection $R$.
The splitting generated by $P$ on vector fields induces a splitting for differential $k$-forms on $\mathcal{M}$. Indeed, given a $k$-form $\alpha$ at some point of $\mathcal{M}$, it can always be written as
$\alpha\,=\,\alpha_{P}^{\parallel}+\alpha_{P}^{\perp}\,,$ (16)
where
$\alpha^{\parallel}_{P}\,:=\,\alpha(P(\,\cdot\,),...,P(\,\cdot\,))\,,$ (17)
and
$\alpha^{\perp}_{P}\,:=\,\alpha-\alpha^{\parallel}_{P}\,.$ (18)
We will refer to $\alpha^{\parallel}_{P}$ and $\alpha^{\perp}_{P}$ as the parallel and transversal components of $\alpha$ associated with the connection $P$.
In the system of local coordinates chosen, if one writes the $k$-form in terms of the basis of differential $1$-forms
$\langle\left\{\,{\rm d}x^{a},\,{\rm d}f^{A}-P^{A}_{a}{\rm d}x^{a}\,=:\,\theta^{A}\,\right\}_{a=1,...,l;A=1,...,r}\rangle$ (19)
dual to the basis of vertical and horizontal vectors associated to $P$, one has
$\alpha\,=\,a_{j_{1}...j_{k}}\rho^{j_{1}}\wedge...\wedge\rho^{j_{k}}\,,$ (20)
where
$\begin{split}\rho^{j_{i}}\,&=\,{\rm d}x^{a}\,,\;\;\;\text{if}\;\;j_{i}=a\,,\\ \rho^{j_{i}}\,&=\,\theta^{A}\,,\;\;\;\text{if}\;\;j_{i}\,=\,A\,.\end{split}$ (21)
In this system of local coordinates, one gets
$\begin{split}\alpha^{\parallel}_{P}\,&=\,\alpha_{A_{1}...A_{k}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k}}\,,\\ \alpha^{\perp}_{P}\,&=\,\sum_{I^{a}_{i}}\alpha_{I^{a}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}x^{a}}_{i}\wedge...\wedge\rho^{j_{k}}\,,\end{split}$ (22)
where $I^{a}_{i}$ is a multi-index $j_{1}...j_{k}$ where the $i$-th component is $a$ and the sum is taken over all the multi-indices of this type. We will denote by $\mathscr{J}$ the set of these multi-indices. As for vertical and horizontal tangent vectors, it is easy to show that this splitting of $k$-forms is stable under a change of local chart, and, thus, it defines two subbundles of $\Lambda^{k}(\mathcal{M})$. We will denote by
$\Lambda^{k}(\mathcal{M})\,=\,{\Lambda^{k}}^{\parallel}_{P}(\mathcal{M})\oplus{\Lambda^{k}}^{\perp}_{P}(\mathcal{M})\,,$ (23)
the splitting of the fibre bundle $\Lambda^{k}(\mathcal{M})$ into the direct sum of the fibre subbundles ${\Lambda^{k}}^{\parallel}_{P}(\mathcal{M})$ and ${\Lambda^{k}}^{\perp}_{P}(\mathcal{M})$ whose sections are parallel (resp. transversal) differential $k$-forms with respect to $P$. We will denote by
$\left\{\,q^{j},\,p_{A_{1}...A_{k}}\,\right\}_{j=1,...,d;A_{1},...,A_{k}=1,...,r}\,,$ (24)
a system of local coordinates on ${\Lambda^{k}}^{\parallel}_{P}(\mathcal{M})$ and by
$\left\{\,q^{j},\,p_{I^{a}_{i}}\,\right\}_{j=1,...,d;I^{a}_{i}\in\mathscr{J}}\,,$ (25)
a system of local coordinates on ${\Lambda^{k}}^{\perp}_{P}(\mathcal{M})$.
Analogously, by means of the connection $R$, one can construct the splitting
$\alpha\,=\,\alpha^{\parallel}_{R}+\alpha^{\perp}_{R}\,,$ (26)
where
$\alpha^{\parallel}_{R}\,:=\,\alpha(R(\,\cdot\,),...,R(\,\cdot\,))\,,$ (27)
and
$\alpha^{\perp}_{R}\,:=\,\alpha-\alpha^{\parallel}_{R}\,.$ (28)
In the system of local coordinates chosen
$\begin{split}\alpha^{\parallel}_{R}\,&=\,\alpha_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,,\\ \alpha^{\perp}_{R}\,&=\,\sum_{I^{A}_{i}}\alpha_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k}}\,,\end{split}$ (29)
where $I^{A}_{i}$ is a multi-index $j_{1}...j_{k}$ where the $i$-th component is $A$ and the sum is taken over all the multi-indices of this type. We will denote by $\mathscr{I}$ the set of these multi-indices.
We will denote by
$\Lambda^{k}(\mathcal{M})\,=\,{\Lambda^{k}}^{\parallel}_{R}(\mathcal{M})\oplus{\Lambda^{k}}^{\perp}_{R}(\mathcal{M})\,,$ (30)
the splitting of the fibre bundle $\Lambda^{k}(\mathcal{M})$ into the direct sum of the fibre subbundles ${\Lambda^{k}}^{\parallel}_{R}(\mathcal{M})$ and ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$ whose sections are parallel (resp. transversal) differential $k$-forms with respect to $R$. Clearly, if $k>l$, ${\Lambda^{k}}^{\parallel}_{R}(\mathcal{M})$ is empty, and $\Lambda^{k}(\mathcal{M})\,=\,{\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$.
We will denote by
$\left\{\,q^{j},\,p_{a_{1}...a_{k}}\,\right\}_{j=1,...,d;a_{1},...,a_{k}=1,...,r}\,,$ (31)
a system of local coordinates on ${\Lambda^{k}}^{\parallel}_{R}(\mathcal{M})$ and by
$\left\{\,q^{j},\,p_{I^{A}_{i}}\,\right\}_{j=1,...,d;I^{A}_{i}\in\mathscr{I}}\,,$ (32)
a system of local coordinates on ${\Lambda^{k}}^{\perp}_{P}(\mathcal{M})$.
3 The theorem
Here, we provide the statement and the proof of the multisymplectic coisotropic embedding theorem.
3.1 Statement
Theorem 3.1.
Let $(\mathcal{M},\omega)$ be a $k$-pre-multisymplectic manifold. Then, there exists a multisymplectic manifold $\tilde{\mathcal{M}}$, referred to as the multisymplectic thickening of $\mathcal{M}$, and an embedding $\mathfrak{i}\;:\;\;\mathcal{M}\hookrightarrow\tilde{\mathcal{M}}$, such that:
  1. 1.
    $\tilde{\mathcal{M}}$ is equipped with a $k$-multisymplectic form $\tilde{\omega}$.
  2. 2.
    $(\mathcal{M},\omega)$ becomes a $(k-1)$-coisotropic submanifold of (~$(\tilde{\mathcal{M}},\tilde{\omega})$. In particular $\mathfrak{i}^{\star}\tilde{\omega}\,=\,\omega$.
3.2 Proof
Consider a $d$-dimensional $k$-pre-multisymplectic manifold $(\mathcal{M},\,\omega)$. In the system of coordinates adapted to the foliation induced by the kernel of $\omega$ described in Section 2, $\omega$ reads:
$\omega\,=\,\omega_{a_{1}...a_{k}}(x){\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,.$ (33)
Consider the bundle ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ described in Section 2, with the system of local coordinates
$\left\{\,q^{j},\,p_{I^{A}_{i}}\,\right\}_{j=1,...,d;\,I^{A}_{i}\in\mathscr{I}}\,.$ (34)
Consider the $(k-1)$-form on ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ defined by
${\Theta_{0}^{(k-1)}}_{p}(V_{1},...,V_{k-1})\,=\,p(\pi_{\star}V_{1},...,\pi_{\star}V_{k-1})\,,$ (35)
where $p$ has to be understood as a point in ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$ on the left hand side and as a (transversal) $(k-1)$-form on $\mathcal{M}$ on the right hand side, $\pi$ is the projection from ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$ to $\mathcal{M}$, and $V_{j}\in\mathbf{T}_{p}{\Lambda^{k}}^{\perp}_{R}(\mathcal{M})\,,\;\;\;\forall\,\,j=1,...,k-1$. In the system of local coordinates chosen
${\Theta_{0}^{(k-1)}}\,=\,\sum_{I^{A}_{i}}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge{\rm d}\rho^{j_{k-1}}\,.$ (36)
Consider on ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ the following $k$-form
$\begin{split}\tilde{\omega}\,:=\,\pi^{\star}\omega+{\rm d}\Theta^{(k-1)}_{0}\,.\end{split}$ (37)
In the system of local coordinates chosen
$\begin{split}\tilde{\omega}\,=\,&\omega_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}+\\ +&\sum_{I^{A}_{i}}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i},\,j}(-1)^{j+1}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,.\end{split}$ (38)
The $k$-form $\tilde{\omega}$ is closed by construction.
To prove whether it is also non-degenerate, we will consider its contraction along the following basis of vector fields on $\tilde{\mathcal{M}}$:
$\left\{\,\frac{\partial}{\partial f^{A}},\,W_{a},\,\frac{\partial}{\partial p^{I^{A}_{i}}}\,\right\}_{a=1,...,l;\,A=1,...,r;\,I^{A}_{i}\in\mathscr{I}}\,.$ (39)
Via a direct computation one shows
$i_{\frac{\partial}{\partial f^{A}}}\tilde{\omega}\,=\,$ $\sum_{I^{A}_{i}}(-1)^{i}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{\theta^{A}}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$ (40)
$+$ $\sum_{I^{A}_{i},\,j}(-1)^{j+i}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\widehat{\theta^{A}}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
-- $\sum_{I^{A}_{i},\,j}p_{I^{A}_{i}}\,{\rm d}_{V}\theta^{B}\left(\,\frac{\partial}{\partial f^{A}}\,\right)\,\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{{\rm d}\theta^{B}}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$
$i_{W_{a}}\tilde{\omega}\,=\,$ $\omega_{a_{1}...a...a_{k}}\,{\rm d}x^{a_{1}}\wedge...\wedge\widehat{{\rm d}x^{a}}\wedge...\wedge{\rm d}x^{a_{k}}+$ (41)
$+$ $\sum_{I^{A}_{i}}(-1)^{j}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}x^{a}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
$+$ $\sum_{I^{A}_{i},\,j}(-1)^{i+l}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\widehat{{\rm d}x^{l}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
-- $\sum_{I^{A}_{i},\,j}{\rm d}_{H}\theta^{B}(W^{a})\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{{\rm d}\theta^{B}}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$
$i_{\frac{\partial}{\partial p_{I^{A}_{i}}}}\tilde{\omega}\,=\,$ $\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$ (42)
where ${\rm d}_{V}\theta^{A}\,=\,{\rm d}\theta^{A}(P(\,\cdot\,),\,P(\,\cdot\,))$ and ${\rm d}_{H}\theta^{A}\,=\,{\rm d}\theta^{A}(R(\,\cdot\,),\,R(\,\cdot\,))$. Since none of the right-hand sides of the equations above vanish, one concludes that $\tilde{\omega}$ is non-degenerate.
Remark 3.2.
[…]
Note that the departing $k$-pre-multisymplectic manifold $\mathcal{M}$ can be embedded into $\Lambda^{k-1}(\mathcal{M})$ as the zero-section $\sigma_{0}\,=:\,\mathfrak{i}$ of the projection $\pi$. A straightforward computation also shows that
$\mathfrak{i}^{\star}\tilde{\omega}\,=\,\omega\,.$ (43)
To show that $\mathcal{M}$ is coisotropic inside $\tilde{\mathcal{M}}$, let us consider a vector field tangent to $\mathcal{M}$
$W\,=\,W^{j}(q)\frac{\partial}{\partial q^{j}}\,,$ (44)
and a vector field tangent to $\tilde{\mathcal{M}}$
$V\,=\,V^{j}(q,\,p)\frac{\partial}{\partial q^{j}}+V^{I^{A}_{i}}(q,\,p)\frac{\partial}{\partial p_{I^{A}_{i}}}\,,$ (45)
and let us compute
[…]
A direct computation shows that
$\begin{split}i_{V\wedge W}\tilde{\omega}\,=\,&\omega_{a_{1}...a_{s}...a_{l}...a_{k}}\,(-1)^{j+k+1}\,V^{s}W^{l}\,{\rm d}x^{a_{1}}\wedge...\wedge\widehat{{\rm d}x^{a_{s}}}\wedge...\wedge\widehat{{\rm d}x^{a_{l}}}\wedge...\wedge{\rm d}x^{a_{k}}+\\ +&\sum_{I^{A}_{i}}(-1)^{l+1}V^{I^{A}_{i}}W^{l}\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i}}(-1)^{s+l}V^{s}W^{l}\,{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i},\,j}(-1)^{j+1+l+s}V^{s}W^{l}\,p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,.\end{split}$ (47)
The term involving
${\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}$ (48)
is independent on the others. Thus, imposing $i_{V\wedge W}\tilde{\omega}$ to vanish for all $W\in\mathscr{X}(\mathcal{M})$ forces all the $V^{j}$ to be zero. Imposing this condition into the other terms leads to the fact that also all the $V^{I^{A}_{i}}$ vanish. Consequently, $\mathbf{T}\mathcal{M}^{\perp,1}$ is empty, and, thus, $\mathcal{M}$ is trivially $1$-coisotropic inside $\tilde{\mathcal{M}}$.
The arguments above can be iterated up to the order $(k-1)$ to show that $\mathcal{M}$ is $(k-1)$-coisotropic inside $\tilde{\mathcal{M}}$.
Conclusions and Future Plans
This paper has established a coisotropic embedding theorem for pre-multisymplectic manifolds. We constructed a suitable multisymplectic thickening of any given pre-multisymplectic manifold such that the original manifold can be embedded as a coisotropic submanifold. This embedding facilitates a more accessible finite-dimensional approach to regularization techniques, which traditionally require handling infinite-dimensional spaces.
Building on these results, several directions for future research emerge.
Applications to Field Theories
We will apply the coisotropic embedding theorem proved to Field Theories, particularly in constructing a Poisson bracket on the space of solutions, as we already did in [CDI+24, CDI+22b, CDI+22c]. We will study the relation between the two approaches and the potential advantages of the one based on the multisymplectic coisotropic embedding theorem proved in this manuscript.
Application to the Integrability of Implicit Partial Differential Equations
Implicit partial differential equations can often be formulated using pre-multisymplectic structures. The regularization method introduced in this paper allows for the definition of a multisymplectic structure on an extended manifold, enabling (under some conditions that have to be studied) the formulation of an explicit PDE. Future work will investigate:
  • The correspondence between solutions of the original implicit PDE and the explicit PDE on the extended manifold.
  • The conditions under which the integrability of the explicit PDE ensures the integrability of the original system.
  • The development of tools to apply this framework to specific classes of implicit PDEs, thereby improving the understanding of their geometric and analytic properties.
Extending Local Results to Global Contexts
Several results proven by the author, including those in [CDI+22a, Sch24b, Sch24a], rely on a local formulation of Field Theories as pre-symplectic Hamiltonian systems. Using the coisotropic embedding theorem, we aim to extend these results to global contexts by avoiding passing through the above-mentioned pre-symplectic Hamiltonian system, which is only locally defined close to some Cauchy hypersurface.